OpenStudy (dls):

A 5L cylinder contained 10 moles of oxyegen gas at 300.Due to sudden leakage,all the gas escaped into atm and cylinder got empty,If Atm p=1atm calculate WD by gas

OpenStudy (dls):

WD=-p delta V delta V=Vf-Vi Vf=0 so delta V =0 so WD=0 o_O

OpenStudy (shubhamsrg):

vfinal = 0 ? o.O

OpenStudy (dls):

all the gas escaped into atm and cylinder got empty :o

OpenStudy (anonymous):

v = nrt/p

OpenStudy (shubhamsrg):

Vfinal = nrT/p

OpenStudy (shubhamsrg):

yep, exactly

OpenStudy (anonymous):

w= -p(change in volume)

OpenStudy (dls):

wtf why

OpenStudy (dls):

Vf not 0 why?

OpenStudy (anonymous):

v = nrt/p v= 10*8.314*300/1 = 24942 J w= -p(change in volume) w=-1*24942= -24942 or 24.9 kj this is a different version of your question

OpenStudy (shubhamsrg):

we are calculating work done by gas. v initial = 5 (volume occupied by gas in cylinder) vfinal = nrt/p (volume occupied by gas in air)

OpenStudy (dls):

oh :o thanks!

OpenStudy (shubhamsrg):

(Y)

OpenStudy (anonymous):

good job @shubhamsrg you explain it really easily and simply :)

OpenStudy (dls):

<3

OpenStudy (shubhamsrg):

hahahahha thank you jesus! B|